OpenStudy (anonymous):

Which of the following classes has the largest ratio of boys to girls? class with 8 boys and 8 girls class with 9 boys and 3 girls class with 12 boys and 6 girls class with 6 boys and 12 girls

OpenStudy (johnweldon1993):

Well lets see Class 1) 8 boys to 8 girl = 1:1 ratio Class 2) 9 boys to 3 girls = 3:1 ratio Class 3) 12 boys to 6 girls = 2:1 ratio Class 4) 6 boys to 12 girls = 1:2 ratio so Class 1) For every 1 boy there is 1 girl Class 2) For every 3 boys there is 1 girl (pretty big ratio) Class 3) For every 2 boys there is 1 girl Class 4) For every 1 boy there are 2 girls (small ratio) So I would go with class 2...make sense?

OpenStudy (anonymous):

tenks

OpenStudy (anonymous):

\[\frac{ 8 }{8 }=1\] \[\frac{ 9 }{ 3}=3\] \[\frac{ 12 }{ 6 }=2\] \[\frac{ 6 }{ 12 }=\frac{ 1 }{ 2}\] class 2 has largest ratio